OpenStudy (anonymous):

The vertices of a rectangle are (-3,-3),(-9,-3),(-9,-7),(-3,-7). Draw the rectangle and find its perimeter and area.

OpenStudy (polaris_s0i):

you can do the drawing part, here is how to calculate the length of each of the sides. \[ s^2 = (\Delta x)^2 + (\Delta y)^2 \] \[ s^2 = (x_2 - x_1)^2 + (y_2-y_1)^2 \] \[ s^2 = ((-9) - (-3))^2 + ((-3)-(-3))^2 \] \[ s^2 = (-6)^2 + \cancel{(0)^2} \] \[ \sqrt{s^2} = \sqrt{(-6)^2} \] \[ s = \sqrt{36} \] \[ s = 6 \]

OpenStudy (mertsj):